OpenStudy (anonymous):

Eliminate the parameter. x = 3 cos t, y = 3 sin t

OpenStudy (ipwnbunnies):

Use this trig identity: \[\cos^2 x + \sin^2 x = 1\]

OpenStudy (anonymous):

i dont remember how.....

OpenStudy (ipwnbunnies):

x = 3 cos t; x^2 = 9 cos^2 Do the same for the 'y' side. Then, apply the identity.

OpenStudy (anonymous):

y=3 sin t;y^2=9 sin ^2 9cos^2+9sin^2=1??

OpenStudy (ipwnbunnies):

Nopeee. x^2 + y^2 = 9 cos^2 t + 9 sin^2 = ? 9(cos^2 t + sin^2 t) = 9(1) = 9 x^2 + y^2 = ... = 9

OpenStudy (anonymous):

whats the ...in between the =

OpenStudy (ipwnbunnies):

The 9 cos^2 t + 9 sin^t. I factored out the 9. Then I can apply the trig identity. This parametric equation is for a circle of radius 3.

OpenStudy (ipwnbunnies):

sin^2 t*

OpenStudy (anonymous):

thank you so much.